A hospital employs 250 nurses. In a random sample of 81 of these, the mean number of

hours overtime billed in a particular week was 12.3, and the sample standard deviation was

8.4 hours. Find a 98% confidence interval for the mean number of hours overtime billed

per nurse in this hospital that week?


1
Expert's answer
2021-12-12T17:32:19-0500

Standard error of sample is "s_s = \frac{\sigma}{\sqrt{n}} = \frac{8.4}{\sqrt{81}} \approx 0.93"


From Student's distribution with 11 degrees of freedom and 98% probability get "t_{cr} = 2.718t"

​therefore;"\epsilon = t_{cr} s_s = 2.718 * 0.93= 2.528"

and the confidence interval is (12.3-2.528, 12.3+2.528)= (9.772, 14.828).